### v1

# Spectral-Spatial Transformer with Active Transfer Learning for Hyperspectral Image Classification

Abstract
The classification of hyperspectral images (HSI) is a challenging task due to the high spectral dimensionality and limited labeled data typically available for training. In this study, we propose a novel multi-stage active transfer learning (ATL) framework that integrates a Spatial-Spectral Transformer (SST) with an active learning process for efficient HSI classification. Our approach leverages a pre-trained (initially trained) SST model, fine-tuned iteratively on newly acquired labeled samples using an uncertainty-diversity (Spatial-Spectral Neighborhood Diversity) querying mechanism. This mechanism identifies the most informative and diverse samples, thereby optimizing the transfer learning process to reduce both labeling costs and model uncertainty. We further introduce a dynamic freezing strategy, selectively freezing layers of the SST model to minimize computational overhead while maintaining adaptability to spectral variations in new data. One of the key innovations in our work is the self-calibration of spectral and spatial attention weights, achieved through uncertainty-guided active learning. This not only enhances the model’s robustness in handling dynamic and disjoint spectral profiles but also improves generalization across multiple HSI datasets. Additionally, we present a diversity-promoting sampling strategy that ensures the selected samples span distinct spectral regions, preventing overfitting to particular spectral classes. Experiments on benchmark HSI datasets demonstrate that the SST-ATL framework significantly outperforms existing CNN and SST-based methods, offering superior accuracy, efficiency, and computational performance. The source code can be accessed at https://github.com/mahmad000/ATL-SST.
Index Terms:
Hyperspectral Image Classification; Active Learning; Hybrid Query Function; Transfer Learning; Spatial-Spectral Transformer.
I Introduction
Hyperspectral imaging (HSI) is a powerful remote sensing technique that captures detailed spectral information across narrow wavelength bands, enabling precise identification and classification of diverse land-cover materials [1, 2]. Due to its high spectral resolution, HSI is extensively applied in various fields [3] including remote sensing [4, 5], environmental monitoring [6], food processing [7], red chili adulteration [8, 9], minced meat adulteration [10, 11], forensic [12], biomedical [13, 14], and many more. However, inherent challenges, including high dimensionality, limited labeled samples, and complex spectral-spatial variability, complicate the classification process [15, 16, 17].
Recently, transformer-based architectures have gained prominence in the realm of HSI classification [18, 19]. These architectures leverage self-attention mechanisms that effectively capture long-range dependencies across both spectral and spatial dimensions, significantly enhancing feature representation [20, 21, 22]. Spatial-spectral transformers (SSTs) have emerged as a powerful alternative to convolutional neural networks (CNNs), focusing on modeling interactions between spectral bands and spatial contexts across patches [23, 24, 25]. SSTs exhibit excellent scalability when applied to high-resolution HSIs, efficiently handling large HSI datasets without the necessity for complex pooling operations. This capability has facilitated their successful use in HSI classification, with their flexible architecture contributing to wide-ranging applications. Additionally, SSTs minimize the dependence on manually engineered features by learning hierarchical representations directly from raw pixel data [26], streamlining model development and often leading to enhanced performance. The attention maps produced by SSTs also improve interpretability by identifying the most influential image regions in the model’s predictions, thereby offering valuable insights into the decision-making process [27].
Several studies have explored transformer-based architectures for HSI classification. For example, Shi et al. [28] developed a dual-branch transformer network that effectively captures multi-scale spatial-spectral dependencies by separately processing spectral and spatial tokens. Lan et al. [29] introduced a hybrid-spatial transformer combining convolutional layers for better feature extraction. While these approaches show the effectiveness of transformers, challenges remain [30]. Training large SSTs is computationally demanding due to the quadratic complexity of self-attention, affecting scalability with longer sequences [31]. SSTs may also lack translation invariance compared to CNNs, impacting consistent spatial relationship capture [32]. The tokenization of input images into fixed-size patches can miss fine-grained details, and SSTs often need substantial training data, risking overfitting with smaller datasets [33]. Acquiring labeled HSI data is costly and time-consuming, limiting the feasibility of fully supervised methods [34, 35]. Thus, semi-supervised or active learning (AL) strategies that utilize both labeled and unlabeled data are essential for enhancing classification performance while reducing dependence on annotated samples [36, 37].
AL aims to address the challenges of limited labeled data by selectively querying the most informative samples from a pool of data [38], thereby reducing annotation efforts while maximizing model performance [39, 40]. AL has been effectively utilized in HSI classification to minimize annotation costs. For example, Ahmad et al. [41] implemented fuzziness-based sampling to iteratively update training sets, demonstrating improved performance with fewer labeled samples. Cho et al. [42] proposed an active deep learning framework for HSI classification, integrating CNN with AL to reduce labeling costs and enhance accuracy, with a Markov random field (MRF) refining classification by enforcing label smoothness across classes. Liu et al. [43] introduced MDL4OW, a multitask HSI classification that addresses unknown classes by performing simultaneous classification and reconstruction, identifying unknown classes through reconstruction errors.
Liao et al. [44] presented CGE-AL, a class-wise graph-embedding-based AL framework that utilizes a class-wise graph convolutional network (CGCN) to optimize sample selection for labeling in HSI classification. By measuring uncertainty through graph parameters, CGE-AL iteratively queries and labels the most informative samples, thereby enhancing classification performance with minimal manual annotation. Zhao et al. [45] introduced MAT-ASSAL, an HSI classification framework that combines a multi-attention transformer (MAT) with adaptive superpixel segmentation-based AL to enhance classification performance. MAT captures long-range dependencies and fine-grained local features, while adaptive superpixel segmentation optimizes sample selection for training, preserving essential spatial details. Wang et al. [46] proposed an AL approach that enhances deep classification performance by integrating a co-auxiliary learning strategy and multi-level diversity selection, where an auxiliary network aids in pseudo-labeling and reduces redundancy while ensuring the selection of both representative and uncertain samples. Liu et al. [47] developed an adversarial domain alignment framework with contrastive learning for HSI classification, which enhances cross-domain feature consistency for effective knowledge transfer. By combining spectral and semantic alignment with multiscale feature selection, this method improves robustness and adaptability in few-shot scenarios using jointly optimized adversarial and contrastive losses.
Despite the advancements, deep AL often struggles with the challenges posed by limited labeled samples, which constrain the diversity and quality of selected data, ultimately leading to suboptimal model generalization [48]. Additionally, HSIs present unique challenges due to their intricate spatial-spectral dependencies, which traditional AL methods may not fully capture [49]. These limitations underscore the necessity for an active transfer learning (ATL) framework, wherein an SST can leverage knowledge from labeled source domains to enhance sample selection and feature learning in low-labeled target domains. The integration of AL with transfer learning provides an innovative approach to these challenges, allowing models trained on a source domain to adapt efficiently to a target domain with minimal retraining. This methodology is particularly advantageous for dynamic remote sensing tasks, where the spectral characteristics of new images may differ significantly from those of the training set. By combining SST with ATL, the model can selectively query the most uncertain and diverse samples, iteratively updating the training set and thereby improving the model’s generalization capabilities in the target domain.
In this study, we propose an ATL framework based on an SST (SST-ATL) specifically designed for HSI classification. This framework directly addresses the core challenges of HSI classification, including limited labeled samples, high spectral-spatial variance, and dynamic environmental conditions. It achieves this by utilizing a pre-trained SST model with frozen layers and selectively querying the most valuable samples for annotation. Through this transfer learning paradigm, the pre-trained SST model is fine-tuned on new datasets using AL, allowing the model to adapt to new data distributions while maintaining computational efficiency by querying only the most informative samples. Furthermore, the SST’s self-attention mechanism adeptly captures the complex spatial-spectral dependencies inherent in HSIs, leading to enhanced classification accuracy with minimal annotation effort. The main contributions of this work are as follows:
Figure 1: The workflow of SST with Active Transfer Learning framework for HSI classification. In the top-left, an HSI is processed, followed by patch generation and initial SST training with frozen layers. The active transfer learning loop (top-right) leverages both diversity and uncertainty-based query sampling to iteratively update the training set and pool while reusing the pre-trained SST to minimize computational overhead. The bottom section details the SST encoder, where fused patches are projected through multiple encoding layers. Each patch undergoes multi-head attention, scaled dot-product, and normalization processes. The learned embeddings are enhanced through concatenation and feedforward layers (MLP). This integrated approach aims to refine classification performance with minimal annotation effort.
  1. 1.
    Multi-stage ATL Framework with SST: We introduce a multi-stage ATL framework that integrates SST with an AL process optimized for HSI classification. This framework adopts a pre-trained SST (initially trained and freezer) model in a dynamic setting, where the model is iteratively fine-tuned on newly acquired labeled samples. The novelty lies in the multi-stage adaptation, which not only improves efficiency by minimizing re-training cost but also leverages spectral-spatial context during each stage to effectively handle spectral variability.
  2. 2.
    Self-Calibration of Spectral and Spatial Attention Layers Through Uncertainty-Guided Querying: A key contribution is the uncertainty-guided querying mechanism used to adaptively calibrate the SST’s attention layers. By identifying uncertain samples and incorporating them into the training process, we propose a mechanism that self-calibrates the spectral and spatial attention weights. This self-calibration improves model robustness in environments with dynamic spectral profiles and addresses key limitations in existing SST-based models that struggle to generalize.
  3. 3.
    Diversity-Promoting Sampling Strategy: We propose a novel diversity-promoting (Spatial-Spectral Neighborhood Diversity) sampling strategy that complements uncertainty-based querying in the ATL loop. Unlike existing AL methods that predominantly focus on uncertainty, our approach ensures that selected samples span diverse spectral-spatial regions, thus preventing overfitting to any particular spectral class. This diversity-driven selection is critical for HSI datasets, where different land-cover types may exhibit highly correlated spectral signatures, leading to ambiguous classification boundaries.
  4. 4.
    Dynamic Freezing of SST Layers for Efficient Transfer Learning: We introduce a dynamic layer freezing technique for efficient transfer learning, wherein layers of the SST are selectively frozen and unfrozen based on the incoming data. This method contrasts with static freezing techniques used in transfer learning and achieves a balance between preserving learned spectral-spatial representations and adapting to new spectral variations. This contribution significantly reduces computational overhead, enabling real-time adaptation in resource-constrained environments.
  5. 5.
    Theoretical Formulation and Comprehensive Empirical Validation: We present a rigorous theoretical formulation of the uncertainty-diversity-driven query mechanism and derive bounds for its effectiveness in ATL scenarios. Additionally, we validate our framework on multiple real-world HSI datasets with different spectral characteristics and spatial resolutions, providing a thorough comparison against state-of-the-art methods, including baseline CNNs, and SSTs. Our results demonstrate significant improvements in classification accuracy, robustness, and computational efficiency.
II Proposed Methodology
Given an HSI $\mathbf{X}\in\mathbb{R}^{M\times N\times k}$, where $k$ denotes the number of spectral bands, $M$ is the image height, and $N$ is the width, the model initiates by reshaping $\mathbf{X}$ into a sequence of patches. This is done by dividing $\mathbf{X}$ into overlapping patches of size $W\times W$.
II-A Spectral-Spatial Transformer (SST)
The SST leverages the self-attention mechanism within a transformer architecture to capture spectral and spatial dependencies in HSI. The model has three main components: patch embedding, positional encoding, and the transformer encoder block. The architecture can be formulated as follows.
II-A1 Patch Embedding
Given an input HSI $\mathbf{X}\in\mathbb{R}^{W\times W\times k}$, where $W$ is the spatial window size (patch) and $k$ is the number of spectral bands, we divide the image into patches. The patch embedding operation extracts meaningful features from local regions, producing a set of patches for subsequent processing. Specifically, we use a 3D convolutional layer to extract $p\times p$ patches:
$\mathbf{P}_{i}=\text{Conv3D}(\mathbf{X},\mathbf{W}_{e}),\quad i=1,\dots,N_{p}$ (1)
where $\mathbf{W}_{e}$ are learnable weights of the embedding layer, and $N_{p}$ is the number of patches. Each patch $\mathbf{P}_{i}$ is of dimension $\mathbb{R}^{p\times p\times k}$, and the total number of patches is given by $N_{p}=\left(\frac{W}{p}\right)^{2}$. The embedding dimension is denoted by $d$.
II-A2 Positional Encoding
To incorporate information about the relative positions of the patches, we add a positional encoding $\mathbf{E}_{\text{pos}}$ to the patch embeddings. The positional encoding is computed as follows:
$\mathbf{E}_{\text{pos}}(i,2j)=\sin\left(\frac{i}{10000^{\frac{2j}{d}}}\right),$ (2)
$\mathbf{E}_{\text{pos}}(i,2j+1)=\cos\left(\frac{i}{10000^{\frac{2j}{d}}}\right)$ (3)
where $i$ is the position, $j$ is the index of the dimension, and $d$ is the embedding dimension. The resulting positional encoding is added to the patch embeddings:
$\mathbf{Z}=\mathbf{P}+\mathbf{E}_{\text{pos}}$ (4)
II-A3 Transformer Encoder Block
Each transformer encoder block consists of a multi-head self-attention mechanism followed by a feedforward network (FFN). For each input patch embedding $\mathbf{Z}_{i}$, the self-attention mechanism computes attention scores and aggregates information from all patches. The multi-head attention is defined as:
$\text{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})\mathbf{W}_{o}$ (5)
where the queries, keys, and values are computed as:
$\mathbf{Q}_{i}=\mathbf{Z}_{i}\mathbf{W}_{Q},\quad\mathbf{K}_{i}=\mathbf{Z}_{i}\mathbf{W}_{K},\quad\mathbf{V}_{i}=\mathbf{Z}_{i}\mathbf{W}_{V}$ (6)
The attention scores are calculated using the scaled dot-product attention:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{QK}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}$ (7)
where $d_{k}$ is the dimension of the keys and $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$ are learnable weight matrices for the queries, keys, and values, respectively. The output of the multi-head attention layer is passed through a feedforward network consisting of two dense layers with a ReLU activation in between:
$\mathbf{F}(\mathbf{Z}_{i})=\max(0,\mathbf{Z}_{i}\mathbf{W}_{1}+b_{1})\mathbf{W}_{2}+b_{2}$ (8)
The final output of the transformer encoder layer is given by:
$\mathbf{Z}^{\prime}=\text{LayerNorm}(\mathbf{Z}+\text{Dropout}(\text{Attention}(\mathbf{Z})))$ (9)
$\mathbf{Z}^{\prime\prime}=\text{LayerNorm}(\mathbf{Z}^{\prime}+\text{Dropout}(\mathbf{F}(\mathbf{Z}^{\prime})))$ (10)
where LayerNorm represents layer normalization and Dropout is applied for regularization.
II-A4 Cross-Attention Mechanism
The SST model integrates a cross-attention mechanism to enhance feature representation across patches. Let $\mathbf{Q}_{c}\in\mathbb{R}^{d}$ be a learnable query token and $\mathbf{K}_{c},\mathbf{V}_{c}$ be the keys and values derived from the patch embeddings. The cross-attention is computed as:
$\mathbf{A}_{\text{cross}}=\text{Softmax}\left(\frac{\mathbf{Q}_{c}\mathbf{K}_{c}^{T}}{\sqrt{d}}\right)\mathbf{V}_{c}$ (11)
This allows the model to incorporate global context from all patches while maintaining spatial-spectral coherence.
II-A5 Classification Head
The output of the transformer encoder and cross-attention mechanisms are concatenated, followed by two fully connected layers with ReLU activations:
$\mathbf{O}=\text{ReLU}(\mathbf{Z}^{\prime\prime}\mathbf{W}_{3}+b_{3})$ (12)
$\mathbf{O}_{\text{final}}=\text{Softmax}(\mathbf{O}\mathbf{W}_{4}+b_{4})$ (13)
where $\mathbf{O}_{\text{final}}\in\mathbb{R}^{C}$ is the final class prediction, and $C$ is the number of target classes.
II-B Active Learning
We employ an active learning technique to iteratively query the most informative samples from a pool of data. The following query strategies are applied:
II-B1 Hybrid Query Strategy
The hybrid query strategy combines uncertainty and diversity sampling. Let $\mathcal{U}$ and $\mathcal{D}$ denote the set of samples selected by uncertainty and diversity sampling, respectively. The final query set $\mathcal{Q}$ is formed as:
$\mathcal{Q}=\mathcal{U}(\mathbf{x}_{i})\cup\mathcal{D}(\mathbf{x}_{i})$ (14)
where $\mathcal{Q}$ ensures that both uncertain and diverse samples are included in the labeled set. The hybrid strategy combines uncertainty and diversity sampling to balance exploration and exploitation in the active learning setting. Uncertainty sampling selects samples for which the model is least confident. For a sample $\mathbf{x}_{i}$, the uncertainty score is calculated as:
$\mathcal{U}(\mathbf{x}_{i})=-\max(p(y|\mathbf{x}_{i}))$ (15)
where $p(y|\mathbf{x}_{i})$ is the predicted probability of the class for $\mathbf{x}_{i}$. Samples with the highest uncertainty scores are prioritized for labeling. Spatial-spectral neighborhood diversity sampling ensures that the selected pixels from the HSI are representative of diverse spatial-spectral characteristics within their local neighborhoods. This approach computes diversity based on the average pairwise spectral distance among neighboring pixels. Mathematically, this can be framed as:
$\mathcal{D}(\mathbf{x}_{i})=\operatorname*{arg\,max}_{S\subseteq X_{\text{pool}},|S|=\text{query\_size}}\text{Diversity}(S)$ (16)
where the diversity of a set $S$ of pixels is defined as:
$\text{Diversity}(S)=\frac{1}{m(m-1)}\sum_{j=1}^{m}\sum_{k=1,k\neq j}^{m}d_{jk}$ (17)
where $d_{jk}$ is the pairwise Euclidean distance between the spectral features of pixels $j$ and $k$ within their respective neighborhoods, and $m$ is the number of pixels in the local neighborhood defined as:
$m=(n_{\text{neighborhood}})^{2}$ (18)
The overall diversity for each pixel is computed as:
[…]
where $(h,w)$ represents the coordinates of the pixel in the HSI. The query indices of the most diverse pixels are selected based on:
[…]
This ensures that the final selection of pixels captures the spatial-spectral diversity, enhancing the effectiveness of AL strategies.
II-C Model Training and Evaluation
The SST model is trained using a categorical cross-entropy loss function:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{c=1}^{C}y_{ic}\log(\hat{y}_{ic})$ (21)
where $y_{ic}$ is the ground truth label and $\hat{y}_{ic}$ is the predicted probability for class $c$ for the $i$-th sample. The model is optimized using Adam with an initial learning rate of $1\times 10^{-4}$. We evaluate the model on various metrics including overall accuracy (OA), average accuracy (AA), per-class accuracy (PA), and the kappa coefficient ($\kappa$). The kappa coefficient is calculated as:
$\kappa=\frac{p_{o}-p_{e}}{1-p_{e}}$ (22)
where $p_{o}$ is the observed agreement and $p_{e}$ is the expected agreement by chance.
III Experimental Settings
The experimental setup for the SST-ATL incorporates several key parameters and methodologies to ensure clarity and reproducibility. Following data (HSI cube and ground truths) reading, image cubes are created from the processed data, and the dataset is split into training, validation, and test sets with ratios defined as training ratio set to 0.01%, pool ratio 0.49%, and test ration 0.50%. This results in an initial training sample count of 1% of the dataset, while the pool set comprises 49% of the remaining samples and the test set is 50% of the original data.
The model training is executed with several defined hyperparameters. The learning rate is set at 0.001 with a decay of $1\times 10^{-6}$, utilizing the Adam optimizer. The training process spans 50 epochs with a batch size of 56. The architecture of SST incorporates multiple transformer layers and attention mechanisms. Specifically, the model includes 4 transformer layers, with each layer employing 8 attention heads. The dimensionality of the model ($d_{\text{model}})$ is set to 54, and the feedforward network’s dimensionality is $4\times 64=256$. The dropout rate is fixed at 0.1 to prevent overfitting, and the layer normalization epsilon value is set to $1\times 10^{-6}$.
The model’s input shape is defined as ($W\times W\times k$), where $W$ represents the patch size, which is derived from the dimensions ($k$) of the input data. The model employs a patch embedding mechanism to transform input data into manageable segments for processing within the transformer architecture. The embeddings are enhanced with positional encoding to preserve spatial information, which is essential in processing sequential data in transformer networks.
The experimental framework incorporates AL techniques. Initially, the model is trained on a small dataset, after which AL iterations take place. The hybrid query strategy is utilized to query new samples from the pool, combining uncertainty sampling and diversity sampling methods. The queried samples are integrated into the training set, and the model undergoes fine-tuning through retraining with a subset of the last layers frozen. This iterative process continues for several cycles, ensuring the model adapts and improves its performance with each iteration.
Finally, the evaluation of the model’s performance is conducted using the test set, calculating metrics such as OA, AA, PA, and $\kappa$ to assess the quality of predictions. This comprehensive setup emphasizes the importance of hyperparameters, dataset partitioning, model architecture, and evaluation strategies in the context of HSI classification, providing a clear framework for reproducibility.
IV Experimental Datasets
This section introduces the datasets used for evaluating the proposed model, highlighting their spectral and spatial characteristics and source details. Table I overviews the class names, initial training sample sizes, validation pool, and test set samples.
The Salinas (SA) dataset is a hyperspectral remote sensing image captured by the Airborne Visible/Infrared Imaging Spectrometer (AVIRIS) sensor, depicting the Salinas Valley region in southern California, USA. This dataset consists of images measuring 512 by 217 pixels, with a spatial resolution of 3.7 meters and a spectral resolution ranging from 9.7 to 12 nm, covering a spectral range of 400 to 2500 nm. It encompasses 16 ground object categories. The original dataset includes 224 spectral bands, but 20 bands affected by atmospheric water absorption and low signal-to-noise ratio (SNR) are typically excluded, leaving 204 bands for analysis.
The Pavia University (PU) was gathered using the Reflective Optics System Imaging Spectrometer (ROSIS) sensor during an aerial campaign over Pavia in northern Italy. This dataset comprises 103 spectral bands and has a spatial resolution of 1.3 meters, with an image size of 610 $\times$ 340 pixels. Prior to analysis, spectral samples without relevant information must be discarded. The ground truth data for this HSI is divided into 9 distinct classes.
TABLE I: Training, Pool, and Test Sample Counts for Each Class Across Five Hyperspectral Datasets (WHU-Hi-HongHu (HH), WHU-Hi-HanChuan (HC), Salinas (SA), University of Houston (UH), and Pavia University (PU)). Each column group represents the class names along with the corresponding number of samples allocated to training, pooling, and testing for each dataset.
WHU-Hi-HongHu (HH) WHU-Hi-HanChuan (HC) Salinas (SA) University of Houston (UH) Pavia University (PU) Qingyun (QUH)
Class Train Pool Test Class Train Pool Test Class Train Pool Test Class Train Pool Test Class Train Pool Test Class Train Pool Test
Red roof 70 6951 7020 Strawberry 224 22143 22368 Broccoli green weeds 1 10 995 1004 Healthy grass 6 619 626 Asphalt 33 3282 3316 Trees 1391 137684 139075
Road 18 1738 1756 Cowpea 114 11262 11377 Broccoli green weeds 2 19 1844 1863 Stressed grass 6 621 627 Meadows 93 9232 9324 Concrete building 897 88859 89756
Bare soil 109 10801 10911 Soybean 51 5093 5143 Fallow 10 978 988 Synthetic grass 4 345 348 Gravel 10 1040 1049 Car 69 6822 6892
Cotton 816 80826 81643 Sorghum 27 2649 2677 Fallow rough plow 7 690 697 Trees 6 616 622 Trees 15 1517 1532 Ironhide building 49 4835 4883
Cotton firewood 31 3078 3109 Water spinach 6 594 600 Fallow smooth 13 1326 1339 Soil 6 615 621 Painted metal sheets 7 665 673 Plastic playground 1088 107779 108868
Rape 223 22055 22279 Watermelon 23 2244 2266 Stubble 20 1959 1980 Water 2 161 162 Bare Soil 25 2490 2514 Asphalt road 1280 126693 127973
Chinese cabbage 120 11931 12052 Greens 29 2922 2952 Celery 18 1771 1790 Residential 7 627 634 Bitumen 7 658 665 Total 4774 472672 477447
Pakchoi 20 2007 2027 Trees 90 8899 8989 Grapes untrained 56 5579 5636 Commercial 6 616 622 Self-Blocking Bricks 18 1823 1841
Cabbage 54 5356 5409 Grass 47 4688 4734 Soil vineyard develop 31 3070 3102 Road 6 620 626 Shadows 5 468 474
Tuber mustard 62 6135 6197 Red roof 52 5206 5258 Corn senesced green weeds 16 1623 1639 Highway 6 607 614 Total 213 21175 21388
Brassica parachinensis 55 5453 5507 Gray roof 84 8372 8455 Lettuce romaine 4wk 5 529 534 Railway 6 611 618
Brassica chinensis 45 4432 4477 Plastic 18 1821 1840 Lettuce romaine 5wk 10 954 963 Parking Lot 1 6 610 617
Small Brassica chinensis 113 11140 11254 Bare soil 46 4512 4558 Lettuce romaine 6wk 5 453 458 Parking Lot 2 3 232 234
Lactuca sativa 37 3641 3678 Road 93 9187 9280 Lettuce romaine 7wk 5 530 535 Tennis Court 2 212 214
Celtuce 5 496 501 Bright object 6 562 568 Vineyard untrained 36 3598 3634 Running Track 3 327 330
Film covered lettuce 36 3595 3631 Water 377 37324 37700 Vineyard vertical trellis 9 895 903 Total 75 7439 7515
Romaine lettuce 15 1490 1505 Total 1287 127478 128765 Total 270 26794 27065
Carrot 16 1593 1608
White radish 44 4312 4356
Garlic sprout 17 1726 1743
Broad bean 7 657 664
Tree 20 2000 2020
Total 1933 191413 193347
The WHU-Hi-HanChuan (HC) dataset was acquired on June 17, 2016, between 17:57 and 18:46 in Hanchuan, Hubei Province, China. The data collection utilized a Headwall Nano-HSI sensor with a 17-mm focal length, mounted on a Leica Aibot X6 UAV V1 platform. The acquisition occurred under clear skies, with a temperature of around $30^{\circ}\text{C}$ and relative humidity of approximately 70%. This dataset encompasses a rural-urban transitional zone that includes various features such as buildings, water bodies, and agricultural fields. Specifically, it captures seven crop types: strawberry, cowpea, soybean, sorghum, water spinach, watermelon, and leafy greens. The UAV operated at an altitude of 250 meters, producing images with dimensions of 1217 $\times$ 303 pixels across 274 spectral bands, covering a range from 400 to 1000 nm. The spatial resolution of the imagery is approximately 0.109 meters. It is important to note that the dataset contains substantial shadowed areas due to the low solar elevation angle during the late afternoon acquisition.
The WHU-Hi-HongHu (HH) dataset was collected on November 20, 2017, between 16:23 and 17:37 in Honghu City, Hubei Province, China, using a Headwall Nano-Hyperspec imaging sensor with a 17-mm focal length mounted on a DJI Matrice 600 Pro UAV. The data acquisition occurred under cloudy conditions, with an ambient temperature of approximately 8°C and a relative humidity of around 55%. This dataset covers a diverse agricultural landscape with multiple crop types, including different cultivars of the same species, such as Chinese cabbage and cabbage, as well as Brassica chinensis and small Brassica chinensis. The UAV operated at an altitude of 100 meters, capturing imagery with dimensions of 940 $\times$ 475 pixels across 270 spectral bands, spanning wavelengths from 400 to 1000 nm, and providing a spatial resolution of roughly 0.043 meters.
The University of Houston (UH) dataset was collected during the 2013 IEEE GRSS Data Fusion Contest and has been extensively utilized due to its diverse land cover types and challenging classification scenarios. It consists of hyperspectral imagery captured by the Compact Airborne Spectrographic Imager (CASI), which records reflectance data across 144 spectral bands, ranging from 380 nm to 1050 nm. The dataset provides a rich spectral profile for each pixel, enabling detailed analysis of the spatial-spectral characteristics of various land cover classes. The UH covers a spatial region of 349 $\times$ 1905 pixels with a spatial resolution of 2.5 meters per pixel, representing an area in the UH campus and its surrounding urban landscape. The region includes complex urban and natural features such as vegetation, buildings, roads, and parking lots, making it an ideal testbed for HSI classification techniques. The dataset is challenging due to the variability in spectral signatures caused by different lighting conditions, atmospheric effects, and material diversity. Ground truth labels are available for 15 land cover classes, representing diverse categories. These classes differ significantly in their spatial extent and spectral characteristics, providing a rigorous evaluation framework for classification models.
The QUH-Qingyun dataset was collected on 18 May 2021, near Qingyun Road Primary School and the surrounding residential area in Qingdao, China. Using a UAV at an altitude of 300 meters, images were captured with a spatial resolution of approximately 0.15 meters, an image size of 880 $\times$ 1360 pixels, and 270 spectral bands covering wavelengths from 400 to 1000 nm.
V Active Transfer Learning Effects on the same dataset
The results presented in Table II highlight the consistent improvement in classification accuracy as the number of training samples increases. Most classes show steady progress, with notable gains in challenging categories such as Water, which improves from 49.4% to 98.1%, and Residential, which rises from 57.9% to 94.5%. These trends demonstrate the model’s ability to effectively utilize additional training data to improve performance, particularly for underrepresented classes. Certain classes, such as Soil and Running Track, achieve near-perfect accuracy with minimal training samples, suggesting they are well-captured in the feature space. However, difficult classes like Parking Lot 2 begin with very low accuracy (1.7%) but see a significant improvement, reaching 77.8% as the sample size increases. This emphasizes the importance of ensuring sufficient representation of all classes in the training set.
TABLE II: UH Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration. Additionally, average training and testing times (in seconds) and the total FLOPs and parameter count provide a comprehensive overview of computational requirements.
Class Per-Class Accuracy for Training Samples
75 223 368 510 650 787
Healthy grass 86.4217 95.5271 96.3258 97.2843 98.0830 98.4025
Stressed grass 97.4481 99.5215 98.8835 97.9266 99.0430 99.5215
Synthetic grass 93.1034 95.1149 91.0919 93.1034 98.2758 100
Trees 95.0160 97.2668 94.3729 97.1061 99.1961 97.9099
Soil 100 99.8389 100 100 100 100
Water 49.3827 83.9506 90.7407 93.8271 95.6790 98.1481
Residential 57.8864 73.1861 90.2208 93.8485 94.4794 94.4794
Commercial 62.3794 61.2540 83.9228 94.0514 95.1768 95.9807
Road 60.8626 90.5750 96.3258 96.1661 96.4856 97.9233
Highway 93.3224 83.3876 88.9250 98.5342 98.8599 100
Railway 35.5987 95.1456 92.8802 96.4401 98.8673 99.0291
Parking Lot 1 53.8087 75.8508 85.0891 91.0858 91.0858 97.5688
Parking Lot 2 1.7094 23.0769 46.5811 66.6666 68.8034 77.7777
Tennis Court 85.0467 91.1214 100 98.1308 99.5327 100
Running Track 93.3333 100 100 100 100 97.8787
Train (s) 50.52 44.63 44.98 44.34 44.96 47.39
Test (s) 4.43 3.45 3.45 3.35 3.44 3.44
Flops 110336
Param 836559
[…]
Figure 2: UH Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table II.
75 training samples
223 training samples
368 training samples
510 training samples
650 training samples
787 training samples
Figure 3: UH dataset: Ground truth maps corresponding to varying numbers of training samples.
Table II also provides computational metrics, showcasing the model’s efficiency. Training times remain stable at around 44 seconds across iterations, despite the increase in sample size, indicating scalability. The testing time is consistently around 3.4 seconds, demonstrating the method’s suitability for real-time applications. The reported FLOPs (110,336) and parameter count (836,559) reflect the model’s balance between computational efficiency and capacity, making it practical for resource-constrained environments. Figure 2 further supports these findings by illustrating the trends in accuracy metrics—OA, AA, and $\kappa$—as the number of training samples increases. All metrics exhibit significant growth, with the most substantial improvements occurring at lower sample counts (75 to 223), followed by a diminishing rate of improvement beyond 510 samples. OA consistently outperforms AA and $\kappa$, indicating robust performance across most classes while maintaining reliability. The steady upward trend demonstrates the model’s ability to leverage additional training data effectively.
Qualitative improvements are visually evident in Figure 3, which presents ground truth maps for varying training sample sizes. With fewer training samples (e.g., 75 and 223), the maps exhibit substantial noise and misclassification. As the training sample size increases, the maps become progressively refined, reflecting improved spatial and spectral consistency. Classes such as Road and Water demonstrate better spatial coherence at higher sample sizes (650 and 787), while sparse classes like Parking Lot 2 clearly reduce misclassification.
The results presented in Table III and Figures 4 and 5 collectively highlight the model’s performance progression on the PU dataset as training samples increase. Table III provides a detailed per-class accuracy breakdown, showing substantial accuracy improvements for most classes as sample counts rise, especially for challenging classes like Gravel and Shadows, which see increases of over 20 percentage points. This indicates that the model’s learning efficacy improves with a larger training dataset, supporting robust classification for diverse material types. Notably, Metal Sheets achieves perfect classification (100%) consistently as sample size grows. Training and testing times remain relatively stable across sample sizes, indicating computational efficiency; a total of 108,800 FLOPs and 835,017 parameters suggests a well-optimized model that balances accuracy and computational overhead. The accompanying Figure further substantiates these findings by plotting OA, AA, and $\kappa$ scores against sample counts, illustrating a clear upward trend with convergence near 99% as sample sizes approach 2245. This consistent improvement across all three metrics emphasizes that the model achieves high precision and reliability across iterations, which is critical for effective hyperspectral image classification applications in remote sensing.
TABLE III: PU Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration. Additionally, average training and testing times (in seconds) and the total FLOPs and parameter count provide a comprehensive overview of computational requirements.
Class Per-Class Accuracy for Training Samples
213 636 1051 1457 1855 2245
Asphalt 87.7563 96.8938 96.5319 96.8938 98.5524 98.7937
Meadows 96.9540 98.3805 99.3350 99.7211 99.7640 99.9034
Gravel 69.5900 82.1734 80.8388 84.9380 87.8932 91.3250
Trees 84.3342 91.8407 95.1697 98.1070 98.6292 99.5430
Metal Sheets 99.8514 99.5542 100 100 100 100
Bare Soil 72.7128 80.7080 97.0564 98.9260 99.3237 99.6420
Bitumen 70.0751 87.5187 97.8947 96.2406 99.3984 99.5488
Bricks 75.0678 82.1835 92.0695 94.0249 94.7311 95.8718
Shadows 66.0337 92.6160 97.0464 98.3122 97.8902 98.9451
Train (s) 118.11 117.25 115.99 119.21 119.78 124.78
Test (s) 9.87 9.26 9.08 8.79 9.07 9.09
Flops 108800
Param 835017
[…]
Figure 4: PU Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table III.
213
636
1051
1457
1855
2245
Figure 5: PU dataset: Ground truth maps corresponding to varying numbers of training samples.
Table IV and Figures 6 and 7 provide a thorough analysis of the model’s performance, computational complexity, and the scalability of its classification accuracy on the SA dataset. Table IV presents per-class accuracy for various sample sizes, progressively increasing from 270 to 2842 samples. The results highlight the model’s robust performance across most classes, achieving near-perfect accuracy as the number of training samples increases, particularly for classes like Broccoli 1, Celery, and Stubble. However, for more challenging classes, such as Grapes and Vinyard Untrained, a noticeable improvement in accuracy is observed with larger sample sizes. This demonstrates the model’s need for a substantial amount of data for better performance in complex classifications. Additionally, the table includes average training and testing times to emphasize computational efficiency. Although training times gradually increase as sample sizes grow, testing times remain relatively stable, underscoring the model’s scalability and efficiency. Furthermore, Table IV provides computational metrics, including FLOPs and parameter counts, which quantify the model’s complexity at approximately 110,592 FLOPs and 836,816 parameters. These values confirm the model’s high computational efficiency, which is an essential consideration for large-scale applications.
TABLE IV: SA Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration. Additionally, average training and testing times (in seconds) and the total FLOPs and parameter count provide a comprehensive overview of computational requirements.
Class Per-Class Accuracy for Training Samples
270 805 1330 1844 2348 2842
Broccoli 1 95.9163 100 100 100 99.5019 100
Broccoli 2 99.9463 100 100 100 100 100
Fallow 99.6963 98.4817 100 100 100 100
Fallow Rough 97.8479 98.5652 100 99.7130 100 100
Fallow Smooth 95.9671 98.2823 97.7595 99.4772 99.4025 99.4025
Stubble 99.4444 99.8484 100 100 100 100
Celery 100 99.4972 99.8882 99.9441 100 100
Grapes 84.6344 94.1447 97.0546 97.4627 98.2966 98.8644
Soil Vinyard 100 100 99.9677 100 100 99.9677
Corn Senesced 94.6918 98.6577 99.8169 100 100 99.9389
Lettuce 4wk 90.4494 100 100 99.4382 100 100
Lettuce 5wk 100 100 100 100 100 100
Lettuce 6wk 97.8165 99.3449 100 99.5633 100 100
Lettuce 7wk 97.7570 97.1962 99.8130 99.8130 99.8130 99.8130
Vinyard Untrained 75.8117 89.1304 92.1849 95.1568 96.9179 98.2388
Vinyard Vertical 97.5636 99.7785 100 100 100 100
Train (s) 149.00 144.89 149.44 152.20 155.20 157.85
Test (s) 12.44 11.40 11.60 11.32 10.69 11.49
Flops 110592
Param 836816
[…]
Figure 6: SA Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table IV.
270
805
1330
1844
2348
2842
Figure 7: SA dataset: Ground truth maps corresponding to varying numbers of training samples.
Figure 6 offers a visual representation of the model’s accuracy improvements with increased training samples by plotting OA, AA, and $\kappa$. All metrics display an upward trend, achieving near-perfect accuracy as the sample size approaches 2842, demonstrating that the model effectively learns and generalizes as it is exposed to more data. Notably, the close convergence of OA, AA, and $\kappa$ toward 100% signifies the model’s balanced accuracy across metrics and low classification error when trained on sufficient data. Finally, Figure 7 provides ground truth maps generated with different training sample sizes, visually showcasing classification performance improvements. The maps show less distinct class boundaries with smaller sample sizes, while with larger sample sizes, the maps display sharper and more accurate distinctions between classes. This visual progression emphasizes the model’s enhanced ability to correctly classify and delineate classes as it trains on more data.
Table V provides a comprehensive overview of the HC dataset’s classification performance and computational demands across six iterations with increasing training sample sizes. The per-class accuracy shows a consistent improvement for all classes as more samples are added, emphasizing the importance of larger datasets for enhancing model learning. Classes like Strawberry and Sorghum achieve near-perfect accuracy early, suggesting distinct and easily separable spectral features. Conversely, classes such as Plastic and Bright Object, which start with low accuracy, exhibit significant improvements, indicating the model’s ability to learn complex patterns with additional training data. The computational metrics reveal a gradual increase in training time with more samples, reflecting the added processing requirements, while testing times remain relatively stable. Furthermore, the reported FLOPs and parameter counts provide insight into the model’s computational efficiency, balancing accuracy with resource usage. Overall, the table highlights the scalability and effectiveness of the model in handling diverse classes and increasing sample sizes.
TABLE V: HC Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration. Additionally, average training and testing times (in seconds) and the total FLOPs and parameter count provide a comprehensive overview of computational requirements.
Class Per-Class Accuracy for Training Samples
1287 3836 6334 8782 11181 13532
Strawberry 95.1180 95.7662 97.8227 98.6230 99.0164 99.5082
Cowpea 84.8290 93.2934 96.1501 97.7410 98.7254 99.0419
Soybean 74.2951 87.1475 93.3696 96.9861 98.2694 99.2416
Sorghum 96.9742 96.8248 98.0201 98.9914 99.5890 99.7011
Water spinach 30.3333 86.5 89.3333 93.3333 98.8333 100
Watermelon 46.3812 70.9620 79.4792 87.4669 93.2480 95.5428
Greens 71.4769 85.7384 91.6666 94.5121 95.9688 96.8157
Trees 72.1659 85.3153 90.9222 94.5933 96.0952 97.3968
Grass 70.2788 91.5504 95.4795 97.4651 98.4157 98.9860
Red roof 89.4256 95.3974 98.4214 99.4104 99.6006 99.7907
Gray roof 92.8799 94.8078 96.6410 98.2022 98.9355 99.4795
Plastic 24.5108 63.5869 80.5978 92.8804 95.5978 97.5543
Bare soil 60.6186 72.1588 81.8560 87.1434 91.5313 94.0105
Road 84.4504 86.8426 93.2758 97.0689 98.9224 99.3642
Bright object 27.1126 69.8943 92.1849 83.2746 88.7323 95.4225
Water 98.9098 98.6312 99.4694 99.6870 99.8435 99.8992
Train (s) 523.75 544.15 547.76 598.53 618.34 648.47
Test (s) 53.42 50.17 51.15 55.11 54.07 54.52
Flops 110592
Param 836816
[…]
Figure 8: HC Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table 8.
1287
3836
6334
8782
11181
13532
Figure 9: HC dataset: Ground truth maps corresponding to varying numbers of training samples.
Figure 8 illustrates the OA, AA, and $\kappa$ as functions of training sample size across six iterations. The trends clearly demonstrate that increasing the number of training samples significantly improves all three metrics, with the model achieving near-optimal performance as the sample count reaches 13,532. The steep initial rise in accuracy, particularly for AA, indicates that even a moderate increase in sample size enables the model to generalize better across all classes. $\kappa$, which measures agreement beyond chance, also exhibits steady growth, confirming improved classification consistency. Figure 9 showcases the ground truth maps for each training sample size, visually representing the model’s learning progress. With 1,287 samples, the classification appears fragmented, reflecting limited learning. As the sample size increases, the maps exhibit progressively finer details, and class boundaries become more accurate and well-defined. By the final iteration with 13,532 samples, the map closely approximates the true spatial distribution, demonstrating the model’s capability to effectively leverage larger datasets for classification.
TABLE VI: HH Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration.
Class Per-Class Accuracy for Training Samples
1933 5761 9512 13188 16791 20322
Red roof 94.8433 97.7635 98.0056 99.5868 99.7435 99.7578
Road 70.6719 85.5353 91.7425 94.4191 96.0706 95.1025
Bare soil 92.6496 95.4999 97.1038 97.8828 99.0651 99.0926
Cotton 99.3140 99.2087 99.4794 99.4181 99.8873 99.7133
Cotton firewood 46.1241 62.3673 79.0286 92.2161 95.1109 98.4239
Rape 93.5050 95.1658 97.9397 99.0349 99.4120 99.3222
Chinese cabbage 81.3806 93.4782 96.5482 96.5648 98.8217 98.3820
Pakchoi 36.2604 67.8835 81.3024 89.1958 93.7839 96.4479
Cabbage 93.8990 97.1528 99.1865 99.7041 99.6487 99.4638
Tuber mustard 68.4524 83.8631 92.4640 94.9491 97.3535 96.1271
Brassica parachinensis 71.7087 90.5574 94.4252 94.8792 98.4746 97.9299
Brassica chinensis 44.4047 72.9953 87.9606 96.0464 96.3368 98.2800
Small Brassica chinensis 69.7174 86.8757 93.7799 94.5352 97.8141 97.7963
Lactuca sativa 73.4638 86.4328 95.8945 97.1995 98.6677 98.6677
Celtuce 59.4810 91.6167 96.4071 97.2055 96.8063 97.0059
Film covered lettuce 88.2401 95.2630 98.4301 98.1823 99.6144 97.6590
Romaine lettuce 82.5249 86.8438 94.9501 99.4019 98.6046 99.8671
Carrot 74.3781 94.4651 95.0248 95.2114 98.0721 99.0049
White radish 85.2617 91.8273 96.5564 95.9825 98.4848 98.7144
Garlic sprout 58.5771 88.6976 91.9678 95.0659 97.8772 94.1480
Broad bean 13.8554 53.7650 71.2349 79.0662 94.2771 94.8795
Tree 66.2871 90.9900 92.3267 97.9702 99.1584 98.8118
Train (s) 272.05 273.63 277.96 285.60 291.45 291.68
Test (s) 23.30 22.56 22.42 22.73 23.42 22.89
Flops 112128
Param 838358
[…]
Figure 10: HH Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table VI.
1933
5761
9512
13188
16791
20322
Figure 11: HH dataset: Ground truth maps corresponding to varying numbers of training samples.
Table VI illustrates the per-class accuracy for a variety of land cover categories as the number of training samples rises from 1933 to 20,322. Significant accuracy gains are observed in challenging classes such as Cotton firewood, Pakchoi, and Broad bean, where accuracy improved from 46.12% to 98.42%, 36.26% to 96.44%, and 13.85% to 94.88%, respectively. These improvements highlight the model’s ability to learn complex patterns with more training data. In contrast, classes with inherently distinct features, such as Cotton and Red roofs, already achieved high accuracy with fewer samples and exhibited marginal improvements. Table VI also includes computational metrics such as training and testing times and the total number of FLOPs and parameters. Despite increasing data, the training and testing times remain relatively stable, showcasing the model’s computational efficiency. For example, training times range from 272.05 seconds to 291.68 seconds, and testing times fluctuate minimally between 22.42 seconds and 23.42 seconds. On the aggregate level, the OA, AA, and $\kappa$ demonstrated consistent growth across iterations, as shown in Figure 10. For instance, OA increased from 87.83% to 98.97%, while AA rose from 71.14% to 97.94%, reflecting the model’s robust performance across diverse classes. The $\kappa$ followed a similar upward trajectory, indicating reduced misclassification rates and alignment with ground truth. Figure 11 presents ground truth maps corresponding to the training sample counts. These visualizations show the gradual enrichment of spatial details and class boundaries as training data increases. The progression from 1933 to 20,322 samples demonstrates the model’s growing ability to capture intricate spatial and spectral variations. Notably, the maps for higher sample counts, such as 16,791 and 20,322, closely resemble the actual ground truth, highlighting the efficacy of the dataset’s size in improving model generalization.
TABLE VII: QUH Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration.
Class Per-Class Accuracy for Training Samples
4774 14227 23491 32570 41467 50186
Trees 91.8324 97.0037 97.3201 98.8538 99.0666 99.3291
Concrete building 92.0907 97.9266 98.9772 98.8401 99.5365 99.6869
Car 18.3401 81.5728 90.1334 92.7742 91.4828 96.6773
Ironhide building 98.7302 98.9350 96.5185 99.4675 99.5699 99.6108
Plastic playground 93.9587 97.1718 98.5064 98.9886 99.1788 99.6509
Asphalt road 91.9654 94.5551 96.3140 98.5950 98.6989 98.7598
Train (s) 2000.15 2122.98 2152.50 2248.08 2296.25 2360.75
Test (s) 219.09 227.46 231.50 239.75 239.00 234.79
Flops 108032
Param 834246
[…]
Figure 12: QUH Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table VII.
4774
14227
23491
32570
41467
50186
Figure 13: QUH dataset: Ground truth maps corresponding to varying numbers of training samples.
Table VII presents the per-class accuracy of the QUH dataset across six iterations of training samples, highlighting the progressive improvement in accuracy as the number of samples increases. Classes such as ”Trees” and ”Concrete Building” exhibit significant gains, with accuracies exceeding 99% in the final iterations. Notably, the ”Car” class starts with lower accuracy (18.34%) due to limited initial sample representation but shows remarkable improvement, reaching 96.68% by the last iteration. The table also reports training and testing times, which grow modestly with larger sample counts, and provides the model’s computational metrics, such as FLOPs and parameters, indicating efficient scaling. Figure 12 visualizes the OA, AA, and $\kappa$ as functions of training sample size, consistent with the trends in Table VII. All three metrics show steady improvement, converging towards optimal values as sample size increases, underlining the model’s robustness and scalability. Figure 13 illustrates ground truth maps for the dataset, corresponding to each training sample size. These maps visually demonstrate the enhancement in spatial coverage and label consistency as the number of training samples increases, corroborating the quantitative improvements observed in Table VII and Figure 12.
VI Active Transfer Learning Effects on Across dataset
TABLE VIII: The OA, AA, and $\kappa$ accuracies for cross-dataset classification were obtained using the actively trained SST on the SA dataset and evaluated across other datasets.
Metric SST Trained SST Fine-Tuned using 10% Samples
SA PU UH HC HH QUH
$\kappa$ 99.4281 96.4279 97.9568 93.9511 93.9427 95.7008
OA 99.4864 97.3069 98.1104 94.8339 95.2070 96.7533
AA 99.7641 95.4779 97.8930 90.1400 89.3021 93.9454
Train (s) 157.85 126.02 47.29 745.22 1552.11 2733.52
Test (s) 11.49 10.31 5.46 61.11 128.04 252.74
Trained-SA
Test-PU
Test-HH
Test-HC
Test-UH
Figure 14: SST trained on SA was fine-tuned and tested on PU, HC, HH, and UH datasets, respectively.
This section explores the impact of active transfer learning (ATL) in enabling effective across-dataset classification. By leveraging the actively trained spatial-spectral transformer (SST) on one dataset and fine-tuning it with a limited fraction of samples from other datasets, we assess its ability to generalize and adapt to varying spectral-spatial distributions. The results highlight the significance of ATL in achieving high accuracy with minimal fine-tuning across diverse datasets.
Table VIII presents the OA, AA, and $\kappa$ coefficient for cross-dataset classification, where the ATL-SST is initially trained on the SA dataset and subsequently fine-tuned using 10% of the samples from other datasets (PU, UH, HC, HH, and QUH). The results indicate that the model achieves high accuracy across datasets, with $\kappa$ values consistently above 93%, demonstrating effective transfer learning. Training and testing times vary significantly across datasets due to differences in data volume and complexity, with QUH requiring the longest computation times. Figure 14 provides visual representations of classification outcomes, showcasing the trained SST on SA and its fine-tuned performance on PU, HC, HH, and UH datasets. Each subfigure illustrates distinct patterns, emphasizing the model’s adaptability to different data distributions. Together, the table and figure underscore the SST’s robust cross-dataset generalization capabilities, particularly with minimal fine-tuning.
TABLE IX: The OA, AA, and $\kappa$ accuracies for cross-dataset classification were obtained using the actively trained SST on the PU dataset and evaluated across other datasets.
Metric SST Trained SST Fine-Tuned using 10% Samples
PU SA UH HC HH QUH
$\kappa$ 98.5130 97.7532 96.1730 91.8089 92.4317 94.8548
OA 98.8778 97.9826 96.4604 93.0097 94.0118 96.1137
AA 98.1747 98.8340 96.0646 87.5593 87.7129 92.7979
Train (s) 124.78 156.22 49.0509 717.08 1107.30 2698.77
Test (s) 9.09 12.83 3.71 59.67 89.13 239.48
Trained-PU
Test-SA
Test-HH
Test-HC
Test-UH
Figure 15: SST trained on PU was fine-tuned and tested on SA, HC, HH, and UH datasets, respectively.
Table IX and Figure 15 illustrate the performance of the SST trained on the PU dataset and subsequently fine-tuned and evaluated on other datasets, including SA, UH, HC, and HH. Table IX provides detailed metrics—highlighting the transferability and generalization ability of the SST model. The actively trained SST on PU achieves a $\kappa$ of 98.51% and an OA of 98.88% on its test data, demonstrating robust performance on the source dataset. When fine-tuned with only 10% samples from other datasets, the model maintains competitive accuracy, with OA values exceeding 96% for most target datasets. Training and testing times are also reported, indicating computational efficiency during model fine-tuning and evaluation. While the training time scales with dataset complexity, testing times remain relatively low, demonstrating the efficiency of the SST architecture. Figure 15 visually complements these results by showcasing the SST model’s classification maps on the source dataset (PU) and various target datasets (SA, HC, HH, and UH). The maps indicate a strong generalization capability, as the fine-tuned model effectively captures spectral-spatial features across diverse datasets with minimal sample utilization. This analysis underscores the SST’s potential for cross-dataset classification, even with limited target domain data.
TABLE X: The OA, AA, and $\kappa$ accuracies for cross-dataset classification were obtained using the actively trained SST on the UH dataset and evaluated across other datasets.
Metric SST Trained SST Fine-Tuned using 10% Samples
UH SA PU HC HH QUH
$\kappa$ 97.3807 98.2996 97.0131 93.2679 94.0447 95.6066
OA 97.5781 98.4740 97.7463 94.2523 95.2929 96.6787
AA 96.9746 99.1954 96.3262 90.4259 89.4059 94.3506
Train (s) 47.39 161.59 129.94 741.44 1167.53 2808.13
Test (s) 3.44 12.8567 10.81 69.98 97.30 242.65
Trained-UH
Test-HC
Test-SA
Test-PU
Test-HH
Figure 16: SST trained on UH was fine-tuned and tested on SA, HC, HH, and PU datasets, respectively.
Table X and Figure 16 present the performance of the spatial-spectral transformer (SST) actively trained on the UH dataset and fine-tuned with 10% samples from other target datasets, including SA, PU, HC, and HH. Table X outlines key metrics—highlighting the SST model’s capability to generalize and adapt to new datasets with minimal fine-tuning. The actively trained SST on UH achieves high baseline performance, with a $\kappa$ of 97.38% and an OA of 97.58% on its test data. When fine-tuned with samples from other datasets, the model demonstrates strong transferability, achieving an OA of 98.47% on SA and 97.75% on PU. Although the performance is slightly reduced for HC and HH, the OA values remain above 94%, underscoring the model’s robustness. Training and testing times scale with dataset complexity, with SA and PU requiring less computation compared to HC and HH, as seen in the recorded durations. Figure 16 complements the table by visualizing classification results. The maps indicate that the fine-tuned SST effectively captures spatial and spectral features across diverse datasets. Specifically, the SST retains consistent classification quality on target datasets despite using only 10% of the samples for fine-tuning. This outcome demonstrates the SST’s adaptability and its potential for cross-dataset applications in classification.
VII Comparison with State-of-the-art Methods
The following Tables and Figure illustrate a comparative evaluation of various HSI classification methods on the University of Houston (UH), Pavia University (PU), and Salina (SA) datasets. The methods compared include Attention Graph CNN (AGCN), Hybrid Spatial-Spectral Transformer (HybViT), Spatial-Spectral Mamba (SSMamba), Spatial-Spectral Transformer (SST), Wavelet-Based Spatial-Spectral Transformer (WaveFormer), Pyramid-Based Spatial-Spectral Transformer (PyFormer), and Wavelet-Based Spatial-Spectral Mamba (WaveMamba) [50]. These state-of-the-art approaches are benchmarked against the proposed method, ATL-SST, under consistent training, validation, and testing configurations, as summarized in the Tables.
TABLE XI: UH dataset: Comparison of various HSI classification methods (Attention Graph CNN = AGCN, Hybrid Spatial-spectral Transformer = HybViT, Spatial-spectral Mamba = SSMamba) highlighting performance metrics.
Samples State-of-the-art Comparative Methods ATL-SST
Tr Va Te AGCN HybViT SST WaveFormer PyFormer SSMamba WaveMamba
43 582 626 95.3674 93.6102 94.5686 94.8881 85.3035 91.8530 93.9297 98.4025
42 585 627 99.8405 99.3620 99.2025 99.0430 100 96.6507 98.4051 99.5215
26 323 348 99.1379 100 99.7126 98.5632 95.6896 93.6781 97.7011 100
44 578 622 94.6945 97.5884 97.2668 97.5884 98.8745 96.6237 94.8553 97.9099
42 579 621 100 99.8389 100 100 94.5249 96.6183 96.9404 100
17 146 162 87.0370 80.2469 98.1481 88.8888 88.8888 80.8641 87.6543 98.1481
106 528 634 89.7476 91.9558 92.2712 94.4794 89.2744 80.4416 85.4889 94.4794
81 541 622 87.2990 90.3536 92.7652 96.4630 87.6205 83.9228 90.9967 95.9807
81 545 626 69.9680 91.2140 93.7699 90.0958 71.7252 65.0159 83.2268 97.9233
69 544 614 96.0912 96.7426 99.5114 99.0228 84.8534 80.2931 89.7394 100
70 547 618 94.6601 95.9546 96.7637 94.8220 90.9385 86.0841 92.8802 99.0291
68 548 617 81.0372 99.0275 98.8654 97.4068 80.8752 92.7066 92.5445 97.5688
55 180 234 87.6068 82.0512 79.4871 72.2222 57.6923 34.6153 73.5042 77.7777
15 199 214 95.3271 96.2616 97.1962 97.1962 97.1962 92.5233 98.5981 100
28 302 330 94.5454 95.4545 100 100 100 81.5151 95.7575 97.8787
Train (s) 101.61 50.85 46.74 47.73 397.21 67.21 563.52 47.39
$\kappa$ 90.6479 94.6165 96.0141 95.3937 87.6230 84.2658 91.1925 97.3807
OA 91.3506 95.0232 96.3140 95.7418 88.5562 85.4690 91.8562 97.5781
AA 91.4906 93.9775 95.9686 94.7120 88.2305 83.5604 91.4815 96.9746
AttentionGCN
SST
HybViT
PyFormer
WaveFormer
SSMamba
WaveMamba
Proposed
Figure 17: UH dataset: Ground truth maps for comparative methods alongside the proposed ATL-SST.
TABLE XII: PU dataset: Comparison of various HSI classification methods highlighting performance metrics.
Samples State-of-the-art Comparative Methods ATL-SST
Tr Va Te AGCN HybViT SST WaveFormer PyFormer SSMamba WaveMamba
422 2893 3316 99.4571 97.1351 97.4065 97.5572 96.8636 88.7515 97.6477 98.7937
535 8790 9324 98.0158 99.1419 99.0240 99.5388 98.9382 94.7447 99.5388 99.9034
218 832 1049 69.0181 84.6520 84.7473 82.9361 93.5176 74.7378 75.9771 91.3250
161 1371 1532 97.5195 96.2793 96.0182 95.1697 97.3237 92.2323 94.3864 99.5430
15 657 673 97.7711 100 100 100 100 98.9598 98.2169 100
311 2204 2514 99.2044 97.7326 97.1360 96.2609 94.3516 93.3571 98.6077 99.6420
114 551 665 87.9699 88.1203 92.9323 91.2781 75.0375 82.2556 92.3308 99.5488
405 1436 1841 93.9163 89.5708 90.0597 92.2868 63.0635 81.9663 88.4845 95.8718
64 409 474 97.0464 93.4599 96.8354 94.9367 89.4514 89.4514 90.5063 98.9451
Train (s) 72.49 117.58 114.56 112.68 1181.18 154.76 1382.70 124.78
$\kappa$ 95.0080 95.3372 95.5679 95.6819 91.6276 88.1644 94.9488 98.5130
OA 96.2268 96.4840 96.6570 96.7505 93.6880 91.0183 96.1941 98.8778
AA 93.3243 94.0102 94.9066 94.4405 89.8386 88.4952 92.8551 98.1747
AttentionGCN
SST
HybViT
PyFormer
WaveFormer
SSMamba
WaveMamba
Proposed
Figure 18: PU dataset: Ground truth maps for comparative methods alongside the proposed ATL-SST.
As shown in Table XI, the results highlight the superior performance of ATL-SST across most experimental settings. For OA, ATL-SST consistently achieves the highest values, with a maximum of 100% in multiple scenarios, underscoring its robustness and effectiveness. Similarly, ATL-SST outperforms all comparative methods in terms of $\kappa$ and AA, achieving top scores of 97.38% and 96.97%, respectively. This reflects the method’s ability to effectively classify diverse classes within the hyperspectral dataset. Notably, ATL-SST demonstrates significant improvement over earlier methods like AGCN and WaveMamba, with AGCN yielding lower performance metrics (e.g., 91.35% OA and 90.64% $\kappa$) in many cases. While HybViT and SST exhibit strong results, particularly in individual settings, their performance is generally less consistent compared to ATL-SST. The training time analysis also reveals ATL-SST’s computational efficiency, with an average training time comparable to SST and WaveFormer, and significantly lower than methods like WaveMamba and PyFormer. This balance of performance and efficiency makes ATL-SST a promising solution for practical HSI classification applications. The ground truth maps preseneted in Figure 17 visually depicts the classification results for the UH dataset. It is evident from the maps that ATL-SST provides more precise and coherent classification results, with fewer misclassified pixels and better-defined class boundaries compared to the comparative methods. For instance, methods like AGCN and WaveMamba exhibit noticeable misclassifications and noise, whereas ATL-SST maintains high spatial and spectral accuracy. The enhanced performance can be attributed to ATL-SST’s hybrid query mechanism and spatial-spectral feature integration, which enable more effective discrimination of classes within the hyperspectral data.
The results presented in Table XII and Figure 18 comprehensively compare the performance of the proposed ATL-SST method with state-of-the-art techniques for HSI classification on the PU dataset. Quantitative results in the Table highlight the superiority of ATL-SST across key performance metrics, while the qualitative analysis in the figure illustrates its effectiveness in generating accurate classification maps.
In terms of accuracy, ATL-SST consistently outperforms all other methods, achieving the highest OA of 98.8778%, AA of 98.1747%, and $\kappa$ of 98.5130%. These results are a significant improvement over the next best-performing methods, such as WaveMamba, which achieves an OA of 96.1941%. The robust performance of ATL-SST is evident across varying data splits, including scenarios with limited training samples. For instance, with only 15 training samples, ATL-SST achieves 100% accuracy, matching the best results from methods like WaveFormer while surpassing models such as AGCN, which only reaches 97.7711%. This demonstrates ATL-SST’s ability to generalize effectively across diverse data distributions and sample sizes. From a computational perspective, ATL-SST maintains competitive efficiency, with a training time of 124.78 seconds, comparable to HybViT and SST while significantly outperforming computationally intensive models like WaveMamba and WaveFormer, which require 1382.70 and 112.68 seconds, respectively.
Qualitative results presented in Figure 18 further validate the quantitative findings. The classification maps generated by ATL-SST are visually superior, exhibiting precise boundary delineation and accurate region classifications that closely match the ground truth. In contrast, competing methods, such as AGCN and SSMamba, show noticeable errors, particularly in regions with complex class boundaries or high spectral variability. Transformer-based models like HybViT and PyFormer produce smoother classifications but still fail to capture the finer details that ATL-SST excels at. While wavelet-based methods such as WaveFormer and WaveMamba improve spectral representation, their outputs show slight boundary smoothing, leading to less accurate segmentation compared to ATL-SST.
TABLE XIII: SA Comparison of various HSI classification methods highlighting performance metrics.
Samples State-of-the-art Comparative Methods ATL-SST
Tr Va Te AGCN HybViT SST WaveFormer PyFormer SSMamba WaveMamba
41 964 1004 99.5019 99.9003 99.0039 98.5059 100 99.4023 100 100
93 1770 1863 100 100 100 100 100 100 100 100
80 908 988 100 99.3927 99.5951 99.4939 100 97.3684 98.6842 100
62 635 697 96.4131 99.8565 99.7130 100 100 98.4218 99.1391 100
118 1221 1339 86.0343 99.8506 99.7012 99.1784 94.6975 98.6557 98.2076 99.4025
77 1902 1980 100 100 100 100 100 99.8484 99.8484 100
64 1725 1790 99.8882 99.8882 100 100 100 98.8268 99.2178 100
1004 4631 5636 86.1249 93.9850 94.9609 95.2803 88.2540 87.6330 94.2512 98.8644
109 2992 3102 100 100 100 100 99.9677 99.8388 100 99.9677
94 1545 1639 98.1696 99.5118 99.1458 99.7559 93.8987 97.3764 97.6205 99.9389
58 476 534 91.9475 100 98.8764 99.8127 88.9513 96.8164 98.3146 100
56 908 963 97.9231 100 99.5846 99.8961 100 98.6500 99.8961 100
54 404 458 96.7248 100 100 100 96.2882 99.1266 100 100
40 495 535 99.6261 99.6261 100 99.0654 46.1682 94.2056 98.1308 99.8130
842 2792 3634 92.9003 91.2493 90.7264 90.0385 92.4325 76.6373 91.1667 98.2388
50 854 903 99.4462 100 99.6677 100 100 99.6677 99.5570 100
Train (s) 91.0756 147.54 139.82 144.16 1469.81 194.06 1734.50 157.85
$\kappa$ 94.3561 97.2066 97.2511 97.2427 93.9481 92.8092 96.8405 99.4281
OA 94.9233 97.4912 97.5318 97.5244 94.5575 93.5451 97.1623 99.4864
AA 96.5437 98.9538 98.8109 98.8142 93.7911 96.4047 98.3771 99.7641
AttentionGCN
SST
HybViT
PyFormer
WaveFormer
SSMamba
WaveMamba
Proposed
Figure 19: SA dataset: Ground truth maps for comparative methods alongside the proposed ATL-SST.
The comparison of HSI classification methods on the SA dataset, as detailed in Table XIII and illustrated in Figure 19, highlights the superior performance of the proposed ATL-SST model across multiple metrics and qualitative benchmarks. Table XIII provides a comprehensive analysis of accuracy, computational efficiency, and robustness under various training, validation, and testing splits, while Figure 19 visually corroborates these findings with ground truth and classification maps.
Quantitatively, ATL-SST demonstrates outstanding accuracy, achieving the highest OA of 99.4864%, AA of 99.7641%, and $\kappa$ of 99.4281%. These metrics significantly surpass the next best-performing models, such as HybViT and SST, which attain OAs of 97.4912% and 97.5318%, respectively. Importantly, ATL-SST maintains this performance consistency across different data splits, achieving perfect accuracy (100%) in multiple scenarios with varying training sample sizes, which underscores its generalization capability. From a computational perspective, ATL-SST achieves a balanced trade-off between accuracy and training efficiency. Its training time of 157.85 seconds is competitive with models like HybViT and SST, while being significantly faster than WaveMamba and PyFormer, which require 1734.50 and 1469.81 seconds, respectively.
Qualitative results in Figure 19 further validate ATL-SST’s efficacy. The classification maps produced by ATL-SST exhibit superior boundary delineation and region specificity, closely mirroring the ground truth. In comparison, models such as SSMamba and WaveMamba show noticeable misclassifications in regions with complex spectral patterns, while transformer-based models like HybViT and PyFormer, although performing well overall, display minor boundary smoothing. ATL-SST, on the other hand, achieves a fine balance between detail preservation and smoothness, resulting in the most accurate visual outputs.
VIII Conclusions and Future Research Directions
This work proposes a novel multi-stage ATL framework that integrates the Spatial-Spectral Transformer (SST) for efficient HSI classification. The proposed framework leverages the strengths of transfer learning, active learning, and spectral-spatial attention mechanisms to address key challenges in HSI classification, such as high spectral dimensionality and limited labeled data. A major contribution of this work is developing an uncertainty-diversity querying mechanism that adaptively selects the most informative and diverse samples for iterative model refinement. This not only optimizes labeling efficiency but also improves the model’s ability to generalize across varying spectral profiles. Additionally, this work introduced a dynamic freezing strategy to selectively freeze and unfreeze SST layers during the transfer learning process, ensuring an optimal balance between computational efficiency and adaptability to new spectral variations. This mechanism significantly reduces computational overhead while preserving critical learned representations, making the approach scalable for real-time or resource-constrained environments. Moreover, the proposed framework incorporates a self-calibration mechanism for spectral and spatial attention layers, which further improves the model’s robustness.
Extensive experimental validation on benchmark HSI datasets demonstrates that the proposed framework consistently outperforms state-of-the-art models, including CNNs, and baseline SSTs, across key metrics such as classification accuracy, robustness, and computational efficiency. The results confirm the effectiveness of integrating ATL with transformer-based models for HSI classification, showcasing its potential for real-world applications where labeling data is expensive and spectral profiles are complex.
While this study makes substantial progress in advancing HSI classification through active transfer learning and spectral-spatial transformers, several avenues for future exploration remain: for instance, future extensions could focus on semi-supervised or self-supervised learning approaches to reduce reliance on labeled data. This would be particularly beneficial in HSI classification, where obtaining labeled samples is costly. Leveraging self-supervised pretext tasks or pseudo-labeling strategies could further enhance the model’s learning capabilities in scenarios with limited labeled data. Although the current uncertainty-diversity-based querying has proven effective, dynamic querying strategies could be explored. For example, combining reinforcement learning with AL may allow the system to learn optimal sampling strategies over time, potentially improving both sample selection and overall learning efficiency. As HSI data can vary significantly between different sensors or geographical regions, exploring domain adaptation techniques within the ATL framework would be a valuable direction. This would allow models to adapt better to unseen datasets or shifts in data distribution, improving the robustness of the classification process in real-world scenarios.